OpenStudy (anonymous):

PLEASE HELP! Express 2sin3theta sin5theta as an algebraic sum: a. -cos8theta + cos2theta b. cos8theta - cos2theta c. -cos8theta - cos2theta d. cos8theta + cos2theta

OpenStudy (anonymous):

I keep posting the textbook for a reason http://finedrafts.com/files/Larson%20PreCal%208th/Larson_Precalculus_8th_txtbk.pdf

OpenStudy (anonymous):

Yea, but for some reason it's not loading :/ I appreciate it though

OpenStudy (unklerhaukus):

\[\sin(u ± v) = \sin u \cos v ± \cos u \sin v\] \[\sin3x=\sin(x +2x) = \sin x \cos 2x + \cos x \sin 2x\]

OpenStudy (anonymous):

wait until your computer completely downloads it :)

OpenStudy (anonymous):

hate when that happens lol

OpenStudy (anonymous):

Thanks @UnkleRhaukus can you please clarify? I got d.

OpenStudy (unklerhaukus):

cos8theta + cos2theta is not right, are a. and d. ment to be the same?

OpenStudy (anonymous):

My fault, a. is negative

OpenStudy (anonymous):

There, fixed it

OpenStudy (unklerhaukus):

which one do you get now?

OpenStudy (anonymous):

a.

OpenStudy (unklerhaukus):

good stuff

OpenStudy (anonymous):

THANKS